OpenStudy (anonymous):

there is a law stating that "The ratio of the width to length for the american flag should be 10 to 19." Is a flag measuring 30 by 59 feet of the correct ratio?

OpenStudy (anonymous):

is 19*3 equal to 59?

OpenStudy (anonymous):

I'd use cross-multiplication on this one to test it out. \[\frac{ 10 }{ 19 } = \frac{ 30 }{ 59 }\] \[10 * 59 = 19 * 30\] \[590 = 570\] Does that seem right? :P

OpenStudy (anonymous):

The reason @completeidiot asks you that is because you can see that the law states the flag is to be 10 by 19. If our new width is 3 times the original width (30 = 10 * 3), then our new length must also be 3 times the original, creating his question, is 19 * 3 equal to 59.